OpenStudy (anonymous):

-5(1-5x)+5(-8x-2)=-4x-8x and show work, please? thank you.

OpenStudy (anonymous):

-5(1-5x)+5(-8x-2)=-4x-8x -5+25x-40x-10=-12x -15x-15=-12x -15=3x x=-5